Use the exact values you enter in previous answer(
Answer #1

A)

Change of entropy of Aluminum:

\Delta s=\frac{dQ}{T}=\frac{1085}{44+273}

\boldsymbol{\Delta s=3.4227\frac{J}{K}}

B)

Change of entropy of Copper:

\Delta s=\frac{-dQ}{T}=\frac{-1085}{54+273}

\boldsymbol{\Delta s=-3.318\frac{J}{K}}

C)

Change of entropy of the universe:

\boldsymbol{\Delta s=3.4227-3.318\rightarrow \Delta s=0.1047\frac{J}{K}}